# RecJPQ: Training Large-Catalogue Sequential Recommenders

### 11 Dec 2023 v1

@@ 3.1. Product Quantisation
Product Quantisation is a well-established vectors compression technique, which has been shown to be successful in approximate nearest neighbour methods (Jégou et al., 2011; Johnson et al., 2021; Ge et al., 2014; Li et al., 2020), information retrieval (Khattab and Zaharia, 2020; Santhanam et al., 2022; Izacard et al., 2020) and recommender systems (Balen and Levy, 2019; Kang et al., 2020; Wang et al., 2022). However, Product Quantisation requires the full embeddings tensor to be trained before the embeddings are compressed. Indeed, the quantisation operation is not differentiable, and therefore the model can not be trained end-to-end. Therefore, it requires first training full (non-quantised) model and only after that apply compression. Some recommendation models (e.g. (Kang et al., 2020)) use differentiable variations of Product Quantisation to allow end-to-end training, but these methods still require training full embeddings alongside the quantised versions.
@@
Overall, Product Quantisation addresses Limitations LABEL:limitation:model_tied (it is model agnostic), LABEL:limitation:not-items (it is applicable for training item embeddings compression), and LABEL:limitation:not-similar (when centroids assigned using clusterisation, similar items will have similar codes). However, it does not address Limitations LABEL:limitation:full_embedding (it requires training full embeddings first),and LABEL:limitation:complex_loss (PQ uses a separate loss function for embeddings reconstruction, which is not aligned with the ranking loss). In the next section, we discuss Joint Product Quantisation, a method that can be adapted to addressing these remaining limitations.
@@ 3.2. Joint Product Quantisation
Zhan et al. recently proposed Joint Product Quantisation (JPQ) (Zhan et al., 2021a), a Product Quantisation-based method developed for dense information retrieval. The main difference with the classic product quantisation is that JPQ generates item codes before training the model. As code assignment is the only non-differentiable operation in Product Quantisation, therefore when the codes are assigned before training the model, the model can be trained end-to-end without training full item embeddings – JPQ essentially replaces the embeddings tensor in the model, where item embeddings are constructed via centroid concatenation, as illustrated in Figure 2. Assuming that the codebook in the figure is a constant, all other parameters can be learned using standard gradient descent. In particular, centroid embeddings in JPQ are learnable model parameters, in the same way that item embeddings are learnable parameters in a conventional learned recommendation model. This means that JPQ does not require special loss function components to learn centroids, as they are learned as part of the overall model training process.
Compared to the original Product Quantisation method, JPQ addresses Limitations LABEL:limitation:full_embedding (it does not require training full item embeddings) and LABEL:limitation:complex_loss (it does not require a specific loss function). However, in contrast to plain PQ, JPQ does not provide a mechanism to assign similar embeddings to similar items and therefore does not address Limitation LABEL:limitation:not-similar. The centroid assignments method proposed in the original JPQ paper is specific for text retrieval (as it relies on the existence of a pre-built index for a text document collection generated using the STAR model (Zhan et al., 2021b)). In the next section, we introduce RecJPQ, an adaptation of JPQ to the sequential recommendation scenario, which does not rely on text-specific datasets and models. Our adaptation of JPQ to sequential recommendation requires careful design of novel centroid assignment strategies. Indeed, to the best of our knowledge, this is the first adaptation of the JPQ method to sequential recommendation.
@@
4. RecJPQ
RecJPQ is a Joint Product Quantisation-based method for training recommendation models with a large catalogue of items. As we discuss in Section 3.2, the method used by JPQ for initial centroids assignments relies on the existence of a pre-built index of documents and, therefore, can not be directly used for recommendation scenarios. Hence, the main difference between RecJPQ and the original JPQ is centroid assignment strategies. In general, RecJPQ can be described as follows:
  1. (1)
    Build the item-code mapping matrix (codebook) using one of the centroid assignment strategies (described in Section 4.1).
  2. (2)
    Initialise the centroid embeddings randomly.
  3. (3)
    Replace the item embedding tensor with the concatenation of centroids associated with each item (as illustrated in Figure 2).
  4. (4)
    Train the model end-to-end using the model’s original training task and loss function (this process also trains the centroid embeddings, so they do not need to be trained separately).
The way RecJPQ builds the codebook before training the main model is also similar to how language models (such as BERT (Devlin et al., 2019)) train a tokenisation algorithm before training the main model. Language models also learn embeddings of sub-word tokens instead of learning embeddings of full words to reduce the model’s size; similarly, RecJPQ learns sub-item centroid embeddings instead of learning embeddings of full items. Below, we describe centroid assignment strategies used by RecJPQ.
4.1. Centroid Assignment Strategies
4.1.1. Random Centroid Assignments
@@
In the most simple scenario, we can assign items to centroids randomly. We compose the item code out of $m$ random integers in this case. RecJPQ with random centroids assignments strategy does not address Limitation LABEL:limitation:not-similar (similar items do not have similar codes). Indeed, with random centroid assignments, RecJPQ becomes similar to other ”random” embeddings compression methods, such as the hashing trick (Weinberger et al., 2010) or Quotient Reminder method (Shi et al., 2020b). The main problem with these methods is that unrelated methods are forced to share parts of their representation, which limits the generalisation ability of the models. However, as we show in Section 5, sometimes random assignments may be beneficial, as the random assignments strategy acts as a form of regularisation. Nevertheless, as the random centroid assignments strategy does not address Limitation LABEL:limitation:not-similar (similar items should have similar representations), we introduce further centroid assignment strategies that can address this limitation in the next sections.
121212
4.1.2. Discrete Truncated SVD
@@
As discussed in Section 3.2, the only limitation not addressed by JPQ is Limitation LABEL:limitation:not-similar (similar items should have similar codes). Random centroid assignments discussed in the previous section do not address this limitation either. Hence, in this section, we design a centroids assignments method that addresses this remaining limitation and assigns similar codes to similar items. Some approaches have used item side information, such as textual data for item representations (Rajput, Shashank et al., 2023); however, we address the more generic classic sequential recommendation scenario (which does not rely on item side information) and hence must infer item similarities from the user’s sequences. To achieve this, we employ the SVD algorithm, which has been shown to achieve good results in learning item representations (Zhou et al., 2012) for recommender systems.
@@
We first compute the matrix of sequence-item interactions $M$, where rows correspond to sequences (users), and columns correspond to items. This matrix’s elements $m_{ij}$ are either 1 if $i^{th}$ contains interactions with item $j$ and 0 otherwise. We then compute the truncated SVD decomposition of matrix $M$ with $m$ latent components: $M\approx U\times\Sigma\times V^{T}$, where $U$ is the matrix of user embeddings, $V$ is the matrix of item embeddings, and $\Sigma$ is the diagonal matrix of largest singular values.
Our initial experiments showed that some items have equal embeddings after performing truncated SVD decomposition. This happens when two items interacted with exactly the same set of users. To ensure that all items have different embeddings, we normalise $V$ using the min-max normalisation range and add a small amount of Gaussian noise: $\hat{v}_{ab}=\frac{v_{ab}-\min\limits_{k}{v_{ak}}}{\max\limits_{k}{v_{ak}}-\min\limits_{k}{v_{ak}}}+\mathcal{N}(0,10^{-5});\forall v_{ab}\in{V}$
@@
The variance of the noise ($10^{-5}$) is negligible compared to the range of possible values of normalised embeddings ($[0..1]$ after min-max normalisation). Therefore it has a very small influence on the position of the items in the embeddings space. However, if two items have exactly the same embeddings after decomposition (e.g. this can happen if two items appear in exactly the same set of sequences), the noise allows us to distinguish these two embeddings. Lastly, the assignment of centroids involves discretising each dimension of the normalised item embeddings into $b$ quantiles so that each quantile contains an approximately equal number of items. We use these bins as centroid assignments for the items. Note that although this method requires computing an $m$-dimensional item embeddings tensor (and there can be hundreds of millions of items), it does not require computing them as part of a deep learning model training on a GPU. Indeed, truncated SVD is a well-studied problem. There are effective algorithms for performing it that do not require modern GPUs (Halko et al., 2011). Moreover, as the method only requires computing $m$-dimensionsional embeddings, the table will be many times smaller than full $d$-dimensional embeddings, so the method requires $\frac{d}{m}$ times less memory to store embeddings. Finally, performing truncated SVD is possible in a distributed manner [^13] For example, using Apache Spark
https://spark.apache.org/docs/latest/mllib-dimensionality-reduction
, which makes it is possible to perform truncated SVD even on very large datasets.
141414
In summary, discrete truncated SVD allows assigning similar codes to similar items; it does not require a GPU for intermediate computations and can be easily performed for very large datasets.
4.1.3. Discrete BPR
@@
Truncated SVD is not the only Matrix Factorisation method that can be used for initial centroid assignments. In particular, we also use the classic BPR approach (Rendle et al., 2009) to obtain coarse item embeddings. The method also learns user embeddings (or, in our case, sequence embeddings) $U$ and item embeddings $V$. The estimate of the relevance of an item $i$ for user $j$ is defined as the dot product of user and item embeddings: $r=u_{j}\cdot v_{i}$. In contrast with truncated SVD, BPR does not directly approximate the user-item interaction matrix. Instead, BPR optimises a pairwise loss function that aims to ensure that positive items are scored higher than negative items: $\mathcal{L}_{BPR}=-\log(\sigma(u_{i}\cdot v_{j^{+}}-u_{i}\cdot v_{j^{-}}))$, where $v_{j^{+}}$ is the embedding of a positive item for the user $u$, $v_{j^{-}}$ is the embedding of a randomly sampled negative item, and $\sigma$ is the logistic sigmoid function.
@@
BPR is a very successful and one of the most cited methods in recommender systems, and therefore we use BPR as an alternative strategy for coarse item embedding learning. The rest of the discrete BPR strategy is the same as in the truncated SVD: we also normalise the learned embeddings using min-max normalisation and add a small amount of Gaussian noise to ensure different embeddings for different items. Similar to truncated SVD, BPR does not require learning on a GPU, and there exist distributed implementations [^15] https://github.com/alfredolainez/bpr-spark that allow for learning item embeddings on very large datasets, so overall it can be used as an alternative to truncated SVD for centroids assignments in RecJPQ.
@@
This concludes the description of the centroid assignment strategies for RecJPQ. We now discuss why RecJPQ may act as a regularisation mechanism and improve the performance on the datasets with many long-tail items.
@@ 4.2. RecJPQ as a Regularisation Mechanism
The interactions with items in recommender systems typically have long tail distribution (Park and Tuzhilin, 2008), meaning that few popular items have the most interactions. In contrast, most items comprise the ”long tail” with few interactions. As the training data for these long-tail items is limited, models suffer from overfitting on long-tail items (Zhang et al., 2021), which causes overall performance degradation.
Goodfellow et al. (Goodfellow et al., 2016, Chapter 7.9) argued that one of the most powerful techniques for preventing overfitting is parameters sharing - a technique where certain parameters of the model are forced to be equal. RecJPQ is a special case of parameter sharing: we force different items to share parts of their embeddings. This prevents the model from learning item embeddings that are too specific to only a few training sequences, as each part of the embedding appears in many other sequences via the sharing mechanism.
In our experiments (see Section 5), we indeed observe that RecJPQ may act as a model regulariser and improve the model’s performance; this is especially apparent in the Gowalla dataset, where the proportion of long-tail items is the largest.
4.3. RecJPQ: Summary
In summary, RecJPQ is a model component that takes the place of the item embeddings tensor in sequential recommender systems. RecJPQ is based on the JPQ method, which is a variation of Product Quantisation in turn. RecJPQ addresses all of the limitations described in Section 2: it is model-agnostic (Limitation LABEL:limitation:model_tied); does not require training full embeddings (Limitation LABEL:limitation:full_embedding); does not modify the backbone model’s loss function (Limitation LABEL:limitation:complex_loss); it is suitable for item embeddings compression (Limitation LABEL:limitation:not-items); it can assign similar codes to similar items with the help of discrete truncated SVD or discrete BPR (Limitation LABEL:limitation:not-similar). Futhermore, we argue that RecJPQ may act as a model regulariser, which is an additional advantage when there are many long-tail items in the catalogue. This concludes the description of RecJPQ. In the next section, we experimentally evaluate RecJPQ and analyse its effects on required memory and on the model performance.
161616
@@ 5. Experiments
Table 3. Salient characteristics of experimental datasets. Long tail items are the percentage of items in the catalogue with less than five interactions.
Dataset Users Iems Interactions Average sequence length Long tail items
@@ 1. RQ1
    What are the effects of centroid assignment strategy in RecJPQ?
@@ 5.1. Experimental Setup
5.1.1. Backbone Models
In our experiments, we use two state-of-the-art Transformer-based sequential recommendation models: BERT4Rec (Sun et al., 2019) - a model that uses a transformer encoder based on BERT (Devlin et al., 2019); and SASRec (Kang and McAuley, 2018) - a model which utilises decoder part of the Transformer (similar to GPT (Radford et al., 2019)). For both models, we use the versions [^17] The code for the paper is available at
https://github.com/asash/RecJPQ
from a recent reproducibility paper (Petrov and Macdonald, 2022b), which provides efficient & effective implementations (using the popular Huggingface transformers library (Wolf et al., 2020))
@@ 5.1.1. Backbone Models
. Additionally, to demonstrate that RecJPQ can be applied to other architectures, we use a GRU (Cho and van Merrienboer, 2014)-based model from (Petrov and Macdonald, 2022a) available in the same repository. This model uses the GRU4Rec (Hidasi and Karatzoglou, 2018) architecture, but a slightly different configuration, e.g. it uses LambdaRank (Burges, 2010) as a loss function, which is shown to be effective (Petrov and Macdonald, 2022a).
5.1.2. Datasets
@@
We experiment with three datasets: (i) MovieLens-1M (denoted ML-1M) (Harper and Konstan, 2015) - this is a movie rating dataset that is one of the most popular benchmarks for recommender systems; (ii) Booking.com (Goldenberg and Levin, 2021) - a multi-destination trips dataset, and (iii) Gowalla (Cho et al., 2011) - a check-in dataset. Following common practice (Sun et al., 2019; Petrov and Macdonald, 2022b), we remove users with less than 5 interactions.
Table 3 lists the salient characteristics of the datasets after preprocessing. As can be seen from the table, the number of items in these datasets varies from relatively small (3416 in MovieLens-1M) to large (1,271,638 in Gowalla) - this allows testing RecJPQ in different settings (RecJPQ is designed for large datasets, and we expect it to compress the model by a larger factor on Gowalla).
181818
5.1.3. Evaluation Protocol
Overall, our evaluation protocol follows the protocol from the recent replicability paper (Petrov and Macdonald, 2022b). We use a leave-one-out data splitting strategy: we hold out the last sequence in each sequence in the test set. Additionally, for 1024 randomly selected users, we hold out the second last action into a validation set, which we use for the early stopping mechanism.
We set the maximum sequence length at 200. If the sequence contains more than 200 interactions, we use 200 latest interactions. If the sequence contains less than 200 interactions, we left-pad it to ensure its length is exactly 200.
To ensure that the models are fully converged, following (Petrov and Macdonald, 2022a), we employ an early stopping mechanism on the NDCG@10 metric: we stop training if the metric is not improved for 200 epochs.
5.1.4. Metrics
The main topic of our research is the trade-off between model size and model effectiveness. For measuring effectiveness, following prior research (Kang and McAuley, 2018; Sun et al., 2019; Petrov and Macdonald, 2022b), we use NDCG@10, and as the model size metric, we use the file size of the model checkpoint. Following recent recommendations (Cañamares and Castells, 2020; Krichene and Rendle, 2022; Dallmann et al., 2021), we measure NDCG without using negative sampling.
5.1.5. Baselines
We deploy an adaptation of Quotient Remainder (Shi et al., 2020b) as a baseline compression approach, applied to each base model – this parameter-free hashing-based approach encodes each item using two hashes: the quotient and the remainder of the division of item id by $\left\lceil\sqrt{|I|}\right\rceil$ where $|I|$ is the catalogue size. Quotient Remainder guarantees that each item has a unique code.
We do not apply post-training embedding quantisation (e.g. float16), nor use other methods from Table 1 as baselines, as they are not suitable for our task: EODRec, LightRec, MDQE and MGQE require training full embeddings (we assume that training full embeddings is not an option for a large catalogue), and PreHash is specific for compressing user embeddings, so is not suitable for item embeddings. However, reducing the model size by decreasing the embedding dimensionality can also be seen as a simple baseline. We analyse models using different embedding sizes in Section 5.2.3.
5.2. Results
5.2.1. RQ1. Effect of centroid assignment strategy.
To analyse the effect of the centroid assignment strategy on model performance/model size tradeoff, we compare the original (base) versions of BERT4Rec, SASRec and GRU with RecJPQ versions trained with Random, discrete truncated SVD and discrete BPR centroid assignment strategies. We do not train GRU and BERT4Rec on Gowalla, as these models do not use negative sampling. Training models on this dataset without negative sampling is not feasible due to the large GPU memory requirement for storing output scores (Petrov and Macdonald, 2022a), while applying negative sampling is a substantial change to the models’ training process that is outside of the scope of this paper.
In all cases, we use 512-dimensional embeddings and the code of length $m=8$ (we experiment with other embedding sizes and lengths of the code in the next section). One exception is the base SASRec model on Gowalla; in this case, we use 128-dimensional item embeddings (item embeddings larger than 128 dimensions consume all available GPU memory when embedding compression techniques are not deployed).
@@ 5.2. Results
Table 4. Impact of RecJPQ with different centroid assignment strategies on model size and effectiveness. Relative Size corresponds to model checkpoint size as the percentage of the base model. ${}^{=}$, ${}^{+}$, and ${}^{-}$ denote significance testing results compared to the base, respectively: indistinguishable ($pvalue>0.05$, Bonferroni multi-test correction), better or worse.
@@
191919
5.2.2. RQ2. Effects of code length $m$ and the embedding size on model performance.

### 18 Dec 2023 v2

@@ 3.1. Product Quantisation
Product Quantisation is a well-established vectors compression technique, which has been shown to be successful in approximate nearest neighbour methods (Jégou et al., 2011; Johnson et al., 2021; Ge et al., 2014; Li et al., 2020), information retrieval (Khattab and Zaharia, 2020; Santhanam et al., 2022; Izacard et al., 2020) and recommender systems (Balen and Levy, 2019; Kang et al., 2020; Wang et al., 2022). However, Product Quantisation requires the full embeddings tensor to be trained before the embeddings are compressed. Indeed, the quantisation operation is not differentiable, and therefore, the model can not be trained end-to-end. Therefore, it requires first training full (non-quantised) model and only then applying compression. Some recommendation models (e.g. (Kang et al., 2020)) use differentiable variations of Product Quantisation to allow end-to-end training, but these methods still require training full embeddings alongside the quantised versions.
@@
Overall, Product Quantisation addresses Limitations LABEL:limitation:model_tied (it is model agnostic), LABEL:limitation:not-items (it is applicable for training item embeddings compression), and LABEL:limitation:not-similar (when sub-item ids are assigned using clusterisation, similar items will have similar codes). However, it does not address Limitations LABEL:limitation:full_embedding (it requires training full embeddings first), and LABEL:limitation:complex_loss (PQ uses a separate loss function for embeddings reconstruction, which is not aligned with ranking loss). In the next section, we discuss Joint Product Quantisation, a method that can be adapted to addressing these remaining limitations.
@@ 3.2. Joint Product Quantisation
Zhan et al. recently proposed Joint Product Quantisation (JPQ) (Zhan et al., 2021a), a Product Quantisation-based method developed for dense information retrieval. The main difference with the classic product quantisation is that JPQ generates item codes before training the model. As code assignment is the only non-differentiable operation in Product Quantisation, therefore, when the codes are assigned before training the model, the model can be trained end-to-end without training full item embeddings – JPQ essentially replaces the embeddings tensor in the model, where item embeddings are constructed via sub-item embeddings concatenation, as illustrated in Figure 2. Assuming that the codebook in the figure is a constant, all other parameters can be learned using standard gradient descent on model’s loss. In particular, this means that JPQ does not require special loss function components to learn sub-item embeddings. 101010 Compared to the original Product Quantisation method, JPQ addresses Limitations LABEL:limitation:full_embedding (it does not require training full item embeddings) and LABEL:limitation:complex_loss (it does not require a specific loss function). However, in contrast to plain PQ, JPQ does not provide a mechanism to assign similar embeddings to similar items and, therefore, does not address Limitation LABEL:limitation:not-similar. The sub-item id assignments method proposed in the original JPQ paper is specific for text retrieval (as it relies on the existence of a pre-built index for a text document collection generated using the STAR model (Zhan et al., 2021b)). In the next section, we introduce RecJPQ, an adaptation of JPQ to the sequential recommendation scenario, which does not rely on text-specific datasets and models. Our adaptation of JPQ to sequential recommendation requires careful design of novel sub-item id assignment strategies. Indeed, to the best of our knowledge, this is the first adaptation of the JPQ method to sequential recommendation.
4. RecJPQ
RecJPQ is a Joint Product Quantisation-based method for training recommendation models with a large catalogue of items. As we discuss in Section 3.2, the method used by JPQ for initial sub-item id assignments relies on the existence of a pre-built index of documents and, therefore, can not be directly used for recommendation scenarios. Hence, the main difference between RecJPQ and the original JPQ is sub-item id assignment strategies. In general, RecJPQ can be described as performing the following steps:
Step 1. Build the item-code mapping matrix (codebook) using one of the sub-item id assignment strategies (see Section 4.1).
Step 2. Initialise the sub-item embeddings randomly.
Step 3. Replace the item embedding tensor with the concatenation of sub-item embeddings associated with each item).
Step 4. Train the model end-to-end using the model’s original training task and loss function (this process also trains the sub-item embeddings, so they do not need to be trained separately).
The way RecJPQ builds the codebook before training the main model is also similar to how language models (such as BERT (Devlin et al., 2019)) train a tokenisation algorithm before training the main model. Language models also learn embeddings of sub-word tokens instead of learning embeddings of full words to reduce the model’s size; similarly, RecJPQ learns sub-item embeddings instead of learning embeddings of full items. Below, we describe sub-item id assignment strategies used by RecJPQ.
@@ 4. RecJPQ
4.1. Sub-Item Id Assignment Strategies
@@
Random Sub-Item Id Assignments. In the most simple scenario, we can assign items to sub-item ids randomly. We compose the item code out of $m$ random integers in this case. RecJPQ with random sub-item id assignments strategy does not address Limitation LABEL:limitation:not-similar (similar items do not have similar codes). Indeed, with random sub-item id assignments, RecJPQ becomes similar to other ”random” embeddings compression methods, such as the hashing trick (Weinberger et al., 2010) or the Quotient Reminder method (Shi et al., 2020b). The main problem with these methods is that unrelated methods are forced to share parts of their representation, which limits the generalisability of the models. However, as we show in Section 5, sometimes random assignments may be beneficial, as the random assignments strategy acts as a form of regularisation. Nevertheless, as the random sub-item id assignments strategy does not address Limitation LABEL:limitation:not-similar (similar items should have similar representations), we introduce further sub-item id assignment strategies that can address this limitation in the next sections.
@@
Discrete Truncated SVD. As discussed in Section 3.2, the only limitation not addressed by JPQ is Limitation LABEL:limitation:not-similar (similar items should have similar codes). Random sub-item id assignments, as discussed above, do not address this limitation either. Hence, in this section, we design a sub-item id assignments method that addresses this remaining limitation and assigns similar codes to similar items. Some approaches have used side information, such as textual data for item representations (Rajput, Shashank et al., 2023); however, we address the more generic classic sequential recommendation scenario and, therefore, must infer item similarities from the user’s sequences. To do this, we employ the SVD algorithm, which has been shown to achieve good results in learning item representations (Zhou et al., 2012) for recommender systems.
@@
We first compute a matrix of sequence-item interactions $M$, where rows correspond to sequences (users), and columns correspond to items. This matrix’s elements $m_{ij}$ are either 1 if $i^{th}$ contains interactions with item $j$ and 0 otherwise. We then compute the truncated SVD decomposition of matrix $M$ with $m$ latent components: $M\approx U\times\Sigma\times V^{T}$, where $U$ is the matrix of user embeddings, $V$ is the matrix of item embeddings, and $\Sigma$ is the diagonal matrix of largest singular values.
@@
Our initial experiments showed that some items have equal embeddings after performing truncated SVD decomposition. This occurs when two items have been interacted with by exactly the same set of users. To ensure that all items have different embeddings, we normalise $V$ using the min-max normalisation range and add a small amount of Gaussian noise: $\hat{v}_{ab}=\frac{v_{ab}-\min_{k}{v_{ak}}}{\max_{k}{v_{ak}}-\min_{k}{v_{ak}}}+\mathcal{N}(0,10^{-5});\forall v_{ab}\in{V}$ The variance of the noise ($10^{-5}$) is negligible compared to the range of possible values of normalised embeddings ($[0..1]$ after min-max normalisation). Therefore it has a very small influence on the position of the items in the embeddings space. However, if two items have exactly the same embeddings after decomposition (e.g. this can happen if two items appear in exactly the same set of sequences), the noise allows us to distinguish these two embeddings. Lastly, the assignment of sub-item id involves discretising each dimension of the normalised item embeddings into $b$ quantiles so that each quantile contains an approximately equal number of items. We use these bins as sub-item id assignments for the items. Note that although this method requires computing an $m$-dimensional item embeddings tensor (and there can be hundreds of millions of items), it does not require computing them as part of a deep learning model training on a GPU. Indeed, truncated SVD is a well-studied problem. There are effective algorithms for performing it that do not require modern GPUs (Halko et al., 2011). Moreover, as the method only requires computing $m$-dimensionsional embeddings, the table will be many times smaller than full $d$-dimensional embeddings, so the method requires $\frac{d}{m}$ times less memory to store embeddings. Finally, performing truncated SVD is possible in a distributed manner (e.g., using Apache Spark (spa, 2023)), which allows truncated SVD to be applied on large datasets. 121212 In summary, discrete truncated SVD allows assigning similar codes to similar items; it does not require a GPU for intermediate computations and can be easily performed for very large datasets.
@@
Discrete BPR. Truncated SVD is not the only Matrix Factorisation method that can be used for initial sub-item id assignments. In particular, we also use the classic BPR approach (Rendle et al., 2009) to obtain coarse item embeddings. The method also learns user embeddings (or, in our case, sequence embeddings) $U$ and item embeddings $V$. The estimate of the relevance of an item $i$ for user $j$ is defined as the dot product of user and item embeddings: $r=u_{j}\cdot v_{i}$. In contrast with truncated SVD, BPR does not directly approximate the user-item interaction matrix. Instead, BPR optimises a pairwise loss function that aims to ensure that positive items are scored higher than negative items: $\mathcal{L}_{BPR}=-\log(\sigma(u_{i}\cdot v_{j^{+}}-u_{i}\cdot v_{j^{-}}))$, where $v_{j^{+}}$ is the embedding of a positive item for the user $u$, $v_{j^{-}}$ is the embedding of a randomly sampled negative item, and $\sigma$ is the sigmoid function.
@@
BPR is one of the most cited methods in recommender systems; therefore, we use BPR as an alternative strategy for coarse item embedding learning. The rest of the discrete BPR strategy is the same as in the truncated SVD: we also normalise the learned embeddings using a min-max normalisation and add a small amount of Gaussian noise to ensure different embeddings for different items are obtained. Similar to truncated SVD, BPR does not require learning on a GPU, and there exist distributed implementations (Luke de Oliveira, 2018) that allow for learning item embeddings on very large datasets, so overall, it can be used as an alternative to truncated SVD for sub-item id assignments in RecJPQ.
@@
This concludes the description of the sub-item id assignment strategies for RecJPQ. We now discuss why RecJPQ may act as a regularisation mechanism and improve the performance on the datasets with many long-tail items.
@@ 4.2. RecJPQ as a Regularisation Mechanism
The interactions with items in recommender systems typically exhibit a long tail distribution (Park and Tuzhilin, 2008), meaning that a few popular items receive the most interactions. In contrast, most items comprise the ”long tail” with few interactions. As the training data for these long-tail items is limited, recommendation models can suffer from overfitting on such items (Zhang et al., 2021), causing overall performance degradation.
Goodfellow et al. (Goodfellow et al., 2016, Ch. 7) argued that one of the most powerful regularisation techniques is parameter sharing – a technique where certain parameters of the model are forced to be equal. RecJPQ is a special case of parameter sharing: we force different items to share parts of their embeddings. This prevents the model from learning embeddings that are too specific to only a few training sequences, as each part of the embedding appears in many other sequences via the sharing mechanism. In our experiments (Section 5), we indeed observe that RecJPQ may act as a model regulariser and improve the model’s performance; this is especially apparent in the Gowalla dataset, where the proportion of long-tail items is the largest.
RecJPQ Summary. To conclude Section 4, we summarise that RecJPQ is a model component that takes the place of the item embeddings tensor in sequential recommender systems. RecJPQ is based on the JPQ method, which is a variation of the PQ method. RecJPQ addresses all of the limitations described in Section 2: it is model-agnostic (Limitation LABEL:limitation:model_tied); does not require training full embeddings (Limitation LABEL:limitation:full_embedding); does not modify the backbone model’s loss function (Limitation LABEL:limitation:complex_loss); it is suitable for item embeddings compression (Limitation LABEL:limitation:not-items); it can assign similar codes to similar items with the help of discrete truncated SVD or discrete BPR (Limitation LABEL:limitation:not-similar). Futhermore, we argue that RecJPQ may act as a model regulariser, which is an additional advantage when there are many long-tail items in the catalogue. This concludes the description of RecJPQ. In the next section, we experimentally evaluate RecJPQ and analyse its effects on required memory and on the model performance.
131313
@@ 5. Experiments
Table 3. Characteristics of the datasets. ”Long tail %” is the percentage of items with less than 5 interactions.
Dataset Users Iems Interactions Avg. length Long tail %
@@ 1. RQ1
    What is the effect of the sub-item id assignment strategy?
@@ 5.1. Experimental Setup
Backbone Models. In our experiments, we use two state-of-the-art Transformer-based recommendation models: BERT4Rec (Sun et al., 2019) – a model that uses a transformer encoder based on BERT (Devlin et al., 2019); and SASRec (Kang and McAuley, 2018) – a model which utilises decoder part of the Transformer (similar to GPT (Radford et al., 2019)). For both models, we use the versions [^14] The code for the paper is available at https://github.com/asash/RecJPQ from our recent reproducibility paper (Petrov and Macdonald, 2022b), which provides efficient & effective implementations (using the Huggingface Transformers library (Wolf et al., 2020)). Additionally, to demonstrate that RecJPQ can be applied to other architectures, we use a GRU (Cho and van Merrienboer, 2014)-based model from (Petrov and Macdonald, 2022a) available in the same repository. This model uses the GRU4Rec (Hidasi and Karatzoglou, 2018) architecture, but a slightly different configuration, e.g. it uses LambdaRank (Burges, 2010) as a loss function, which is shown to be effective (Petrov and Macdonald, 2022a).
@@
Datasets. We experiment with three datasets: (i) MovieLens-1M (denoted ML-1M) (Harper and Konstan, 2015) – this is a movie rating dataset that is one of the most popular benchmarks for recommender systems; (ii) Booking.com (Goldenberg and Levin, 2021) – a multi-destination trips dataset, and (iii) Gowalla (Cho et al., 2011) – a check-in dataset. Following common practice (Sun et al., 2019; Petrov and Macdonald, 2022b), we remove users with less than five interactions. Table 3 lists the salient characteristics of the datasets after preprocessing. As can be seen from the table, the number of items in these datasets varies from relatively small (3416 in MovieLens-1M) to large (1,271,638 in Gowalla) – this allows testing RecJPQ in different settings (RecJPQ is designed for large datasets, and we expect it to compress the model by a larger factor on Gowalla).
@@
Table 4. Impact of RecJPQ with different sub-item id assignment strategies on model size and effectiveness for the MovieLens-1M and Booking.com datasets. Relative Size corresponds to model checkpoint size as the percentage of the base model. ${}^{=}$, ${}^{+}$, and ${}^{-}$ denote significance testing results compared to the base, respectively: indistinguishable ($pvalue>0.05$, Bonferroni multi-test correction), better or worse. Bold denotes the best NDCG@10 in each column.
@@
Evaluation Protocol. Overall, our evaluation protocol follows the protocol from the recent replicability paper (Petrov and Macdonald, 2022b). We use a leave-one-out data splitting strategy: we hold out the last sequence in each sequence in the test set. Additionally, for 1024 randomly selected users, we hold out the second last action into a validation set, which we use for the early stopping mechanism. We set the maximum sequence length at 200. If the sequence contains more than 200 interactions, we use 200 latest interactions. If the sequence contains less than 200 interactions, we left-pad it to ensure its length is exactly 200. To ensure that the models are fully converged, following (Petrov and Macdonald, 2022a), we employ an early stopping mechanism on the NDCG@10 metric: we stop training if the metric is not improved for 200 epochs.
Metrics. Our main focus is the trade-off between model size and model effectiveness. For measuring effectiveness, following prior research (Kang and McAuley, 2018; Sun et al., 2019; Petrov and Macdonald, 2022b), we use NDCG@10, and as the model size metric, we use the file size of the model checkpoint. Following best practices (Cañamares and Castells, 2020; Krichene and Rendle, 2022; Dallmann et al., 2021), we measure NDCG without using negative sampling.
151515
Baselines. We deploy an adaptation of Quotient Remainder (Shi et al., 2020b) as a baseline compression approach, applied to each base model – this parameter-free hashing-based approach encodes each item using two hashes: the quotient and the remainder of the division of item id by $\left\lceil\sqrt{|I|}\right\rceil$ where $|I|$ is the catalogue size. Quotient Remainder guarantees that each item has a unique code.
We do not apply post-training embedding quantisation (e.g. float16), nor use other methods from Table 1 as baselines, as they are not suitable for our task: EODRec, LightRec, MDQE and MGQE require training full embeddings (we assume that training full embeddings is not an option for a large catalogue), and PreHash is specific for compressing user embeddings, so is not suitable for item embeddings. However, reducing the model size by decreasing the embedding dimensionality can also be seen as a simple baseline. We analyse models using different embedding sizes in RQ3.
To analyse the effect of the sub-item id assignment strategy on model performance/model size tradeoff, we compare the original (base) versions of BERT4Rec, SASRec and GRU with RecJPQ versions trained with Random, discrete truncated SVD and discrete BPR sub-item id assignment strategies. We do not train GRU and BERT4Rec on Gowalla, as these models do not use negative sampling. Training models on this dataset without negative sampling is not feasible due to the large GPU memory requirement for storing output scores (Petrov and Macdonald, 2022a), while applying negative sampling is a substantial change to the models’ training process that is outside of the scope of this paper. In all cases, we use 512-dimensional embeddings and the code of length $m=8$ (we experiment with other embedding sizes and lengths of the code in the next section). One exception is for the SASRec base model on Gowalla; in this case, we use 128-dimensional item embeddings (item embeddings larger than 128 dimensions consume all available GPU memory when embedding compression techniques are not deployed).
5.2. Results